If TR= 2+3Q^2, then MR is:


1
Expert's answer
2021-01-15T09:46:56-0500

By the definition, the marginal revenue is equal to change in total revenue divided by change in quantity sold:


"MR=\dfrac{\Delta TR}{\Delta Q}."

Therefore, we can write:


"MR=\dfrac{dTR}{dQ}=\dfrac{d}{dQ}(2+3Q^2)=6Q."

Answer:

"MR=6Q."